true or false: (-5, 3, -2) is a solution to the equation 2x+3y-z=1
Answer by checkley71(8403)

2X+3Y-Z=1
2*-5+3*3-(-2)=1
-10+9+2=1
-10+11=1
1=1 TRUE.